OpenStudy (anonymous):

From the set {1, 3, 5}, use substitution to determine which value of x makes the equation true. 10x + 8 = 38 A. none of these B. 1 C. 3 D. 5

OpenStudy (anonymous):

WIll medal and fan

JoelTheBoss (joel_the_boss):

Do you anything about this?

JoelTheBoss (joel_the_boss):

@trinajoy Im going to take that as a no....

OpenStudy (anonymous):

sorry iwas helping someone

OpenStudy (anonymous):

i thought is would be A.

JoelTheBoss (joel_the_boss):

Ok, For this type of problem all you must do is plug in the number. I will start with 3, 10x+8=38 \[10(3)+8=38\]\[10\times3=30\] \[30+8=38\]\[38=38 \]So it would be true! 3 is the correct choice. :)

JoelTheBoss (joel_the_boss):

Hope I helped. :)

OpenStudy (anonymous):

Thank u would u mind is u can help with a few more :)

OpenStudy (anonymous):

U did Thank u

OpenStudy (anonymous):

i just have a few more case am doing a test of 28 questions and it is do today :( so i don't have much time

JoelTheBoss (joel_the_boss):

tag me :)

OpenStudy (anonymous):

Which value for x makes the sentence true? A. B. C. D.

OpenStudy (anonymous):

it's fractions hold on

OpenStudy (anonymous):

1/2X = 9/10 this is the questions

OpenStudy (anonymous):

A.7/5 B.9/20 C.2/5 D.9/5

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@Joel_the_boss